# VSC-RL: Advancing Autonomous Vision-Language Agents with Variational Subgoal-Conditioned Reinforcement Learning

### v1

@@ Abstract
State-of-the-art (SOTA) reinforcement learning (RL) methods enable the vision-language agents to learn from interactions with the environment without human supervision. However, they struggle with learning inefficiencies in tackling real-world complex sequential decision-making tasks, especially with sparse reward signals and long-horizon dependencies. To effectively address the issue, we introduce Variational Subgoal-Conditioned RL (VSC-RL), which reformulates the vision-language sequential decision-making task as a variational goal-conditioned RL problem, allowing us to leverage advanced optimization methods to enhance learning efficiency. Specifically, VSC-RL optimizes the SubGoal Evidence Lower BOund (SGC-ELBO), which consists of (a) maximizing the subgoal-conditioned return via RL and (b) minimizing the subgoal-conditioned difference with the reference policy. We theoretically demonstrate that SGC-ELBO is equivalent to the original optimization objective, ensuring improved learning efficiency without sacrificing performance guarantees. Additionally, for real-world complex decision-making tasks, VSC-RL leverages the vision-language model to autonomously decompose the goal into feasible subgoals, enabling efficient learning. Across various benchmarks, including challenging real-world mobile device control tasks, VSC-RL significantly outperforms the SOTA vision-language agents, achieving superior performance and remarkable improvement in learning efficiency.
Machine Learning, ICML
@@ 1 Introduction
In recent years, the vision-language model (VLM) has shown its significant ability of multimodal understanding and commonsense reasoning (Yang et al., 2023a; Shen et al., 2024; Hong et al., 2023), and has achieved impressive achievements in various real-world applications, such as visual question answering and visual captioning (Bai et al., 2023; Chen et al., 2024). Therefore, it shows great potential to solve real-world complex decision-making problems (e.g., mobile device control (Toyama et al., 2021) and web control (Zhou et al., 2023)) via building intelligent vision-language agents through the advanced vision-language model (Yang et al., 2023c; Zheng et al., 2024). In the past decade, reinforcement learning (RL) has achieved impressive achievements in various sequential decision-making tasks including board games (Silver et al., 2017) and video games (Berner et al., 2019). Some recent works (Bai et al., 2024; Qi et al., 2024) suggest that training vision-language agents with the RL methods for addressing sequential decision-making tasks.
@@
Overall, based on the specific training paradigm, vision-language agents can be categorized into three main types: prompting-based, imitation-based, and RL-based agents. Directly leveraging VLMs (e.g., Gemini-1.5-Pro (Team et al., 2024) and GPT-4V (OpenAI, 2023)) to capture the critical information from the multimodal content, prompting-based agents aim to generate action via prompting engineering and retrieving techniques (Yang et al., 2023c, b). The performance of the prompting-based agents is usually poor as the weights of these VLMs can not be updated for challenging tasks. Some works (Zhang & Zhang, 2023; Hong et al., 2024) suggest using imitation learning techniques to fine-tune the open-source VLMs on the human demonstration. However, the performance of imitation-based agents is highly dependent on the quality of the demonstration, thus may lead to suboptimal performance or even hinder the model’s ability to solve the novel tasks out of distribution. Recently, the most promising method is the RL-based agent which enables the VLM to address complex sequential decision-making tasks by applying RL approaches in training VLM-based policy (Bai et al., 2024; Qi et al., 2024). However, existing RL-based agents often suffer from the learning efficiency issue in addressing challenging control tasks with sparse reward signals and complicated goals. In real-world complex scenarios, the goal of the task is usually to be accomplished via applying a long sequence of actions, and then the reward signal is only given when the goal is finished at the end of the horizon, fundamentally leading to the learning efficiency issue.
To address the fundamental issues of RL-based agents mentioned above. Goal-conditional RL have shown remarkable potential in resolving complex control tasks with sparse rewards and long-horizon, such as robotic manipulation (Chane-Sane et al., 2021; Kalashnikov et al., 2018) and navigation (Zhu & Zhang, 2021; Zhang et al., 2017). Goal-conditioned RL problem consists of a set of normal RL problems and corresponding goals (Liu et al., 2022a). In this context, the agent needs to make the decision from the current observation to accomplish the given specific objective within the horizon. Some recent works attempt to address this issue via introducing implicit curriculum (Andrychowicz et al., 2017) or explicit subgoals (Dayan & Hinton, 1992; Chane-Sane et al., 2021). However, these existing goal-conditioned RL approaches often fail to learn a proper policy in the real-world complex sequential decision-making task due to the complicated subgoal generation and curriculum design methodologies.
@@
In this paper, we introduce Variational Subgoal-Conditioned RL (VSC-RL), a novel RL-based agent method for enhancing learning efficiency in real-world complex sequential decision-making tasks From the perspective of variational inference, VSC-RL reformulates the control problem as the variational subgoal-conditioned RL problem which can be solved by utilizing extensive optimization techniques. Additionally, VSC-RL utilizes the significant reasoning and planning capabilities of VLM to autonomously decompose the complex goal into feasible subgoals. Given the generated subgoals, VSC-RL optimizes the objective of SubGoal-Conditioned Evidence Lower BOund (SGC-ELBO), thus effectively improving learning efficiency. Specifically, VSC-RL optimizes the SGC-ELBO via (a) maximizing the subgoal-conditioned RL return of the target agent and (b) minimizing the subgoal-conditioned difference via imitating the reference agent. We theoretically demonstrate that SGC-ELBO is the equivalent objective of the original optimization objective, ensuring both improved efficiency and performance guarantees. Empirical results on various benchmarks validate our statement that VSC-RL significantly outperforms SOTA vision-language agents in both sample efficiency and final performance.
@@
In this paper, literature related to vision-language agents and RL methods is discussed in Section 2. We introduce notations related to goal-conditioned RL, variational RL and subgoal generator in Section 3. In Section 4, we illustrate how to formulate the sequential decision-making problem as a variational subgoal-conditioned RL problem and derive the new optimization objective: GC-ELBO which is theoretically equivalent to the original objective, followed by the practical implementation of VSC-RL. In Section 5, the experimental results over various benchmarks exhibit our VSC-RL agent can achieve superior performance compared to existing SOTAs. Overall, the main contributions of this paper are summarized as follows:
@@
  • We propose VSC-RL, a novel variational subgoal-conditioned RL method for enhancing vision-language agents in resolving real-world sequential decision-making problems.
@@
  • We theoretically show that SGC-ELBO, the optimization objective of the VSC-RL is equivalent to the original optimization objective, effectively improving learning efficiency and maintaining the performance guarantee.
@@
  • We experimentally show that VSC-RL-based vision-language agent significantly outperforms various SOTAs in both learning efficiency and final performance on different evaluation benchmarks.
@@ 2 Related Works
2.1 Vision-language Agents for Decision-making
In real-world complex control tasks requiring capacities in reasoning, planning and content understanding, it is necessary to enable agents with the vision-language models. In particular, the vision-language model (VLM) can process and abstract the image and language content for challenging decision-making tasks, especially in mobile device control tasks. Existing vision-language agents can be categorized as prompting-based, imitation-based, and RL-based agents based on the corresponding learning paradigms. Additionally, recent advances explore using VLM as subgoal generator to decompose the complicated goal in long-horizon tasks.
@@ 2.1 Vision-language Agents for Decision-making
Leveraging the inherent reasoning and planning abilities of prosperity VLMs (e.g., Gemini-1.5-Pro (Team et al., 2024) and GPT-4V (OpenAI, 2023)), the prompting-based agent makes decision via prompting engineering and retrieving techniques. For instance, AppAgent (Yang et al., 2023c) first introduces a unified prompting-based agent method to enable the vision-language model to directly interact with mobile applications by providing the prompts with details of actions. Set-of-Marks (Yang et al., 2023b) proposes a new prompting method to enhance the visual grounding ability of VLM. However, the performance of these prompting-based agents is always sensitive to prompts required manually and carefully designed. Therefore, it is challenging for the prompting-based agent to directly output the correct and desired actions to address real-world complex control problems.
@@
The imitation-based agent learns to mimic the expert behaviors by fine-tuning the policy on human demonstration. Recently, Android in the Wild (AitW) (Rawles et al., 2024) establish large-scale datasets of mobile device control tasks, enabling agents to directly learn from human experience. AutoUI (Zhang & Zhang, 2023) and CogAgent (Hong et al., 2024) fine-tune the VLM-based policies with the AitW dataset, remarkably outperforming the prompting-based agent. In order to adapt the fine-tuned agent to the online environment, Filtered BC (Pan et al., 2024b) introduces online imitation mechanisms to learn from successful online experiences. Unfortunately, these methods rely heavily on high-quality human demonstrations and often struggle to generalize to unseen tasks, limiting their application in diverse real-world scenarios.
@@
Different to prompting-based and imitation-based agents, the RL-based agent can autonomously optimize the policy through trial-and-error interactions with environments, without human supervision. DigiRL (Bai et al., 2024) introduces a unified offline-to-online RL framework that enables agents to learn directly from real-time interactions in dynamic environments, improving performance without the need for curated datasets. DistRL (Wang et al., 2024) builds an asynchronous distributed RL system, allowing training multiple agents in parallel across different environments, thus significantly enhancing scalability and convergence speed. However, these RL-based agents still fundamentally suffer from the learning efficiency issue in challenging sequential decision-making tasks with the sparse reward and long-horizon.
@@
Recent works have shown that VLM can enhance the RL method via its remarkable capacities of reasoning, planning, and content understanding. Recent works suggest adopting VLM in reward-shaping for RL. For instance, VLM-RMs (Rocamonde et al., 2023) demonstrate that VLMs can serve as effective reward models for learning complex skills. VLM can also generate the subgoals to guide the learning process for autonomous driving (Pan et al., 2024a) and robot (Yang et al., 2024) tasks. Nonetheless, it is still an open problem how to effectively integrate the VLM-generated subgoals into RL.
To mitigate these above issues, we present VSC-RL which can autonomously decompose the goal into feasible subgoals by advanced VLM, and then efficiently resolve each subgoal from the principle of variational inference.
@@ 2.2 Goal-conditioned and Variational RL
We introduce the common RL methods in addressing the control tasks, including goal-conditioned RL and variational RL.
@@
Sequential decision-making tasks can be viewed as the goal-conditioned RL probelm (Liu et al., 2022a). Based on the current state, the agent aims to find the optimal policy that guides progress toward the given goal for maximizing the return. Hindsight experience replay (Andrychowicz et al., 2017) introduces an implicit curriculum learning method to enhance learning efficiency and robustness. With the perspective of divide-and-conquer, some approaches suggest that guiding the agent with subgoals as intermediate reward signals via imagination (Chane-Sane et al., 2021; Nair & Finn, 2019) and tree-search (Jurgenson et al., 2020; Parascandolo et al., 2020).
@@
The RL problem can be viewed as the variational inference problem (Levine, 2018) which can be resolved by utilizing extensive optimization tools, thus effectively improving the learning efficiency. Applying the expectation-maximization algorithm in the actor-critic method in RL, VIP (Neumann, 2011) presents a unified variational inference framework. MPO (Abdolmaleki et al., 2018a, b) proposes a series of off-policy RL with entropy regulation in the manner of expectation-maximization. VDPO (Wu et al., 2024) and CVPO (Liu et al., 2022b) apply the variational inference techniques in addressing RL problem with delayed signals and safety constraints, respectively.
This paper aims to show how to formulate the control problem as the variational subgoal-conditioned RL problem from the perspective of variational inference which allows us to resolve the complicated control task by utilizing extensive optimization tools.
@@ 3 Preliminaries
We formulate the RL problem as the finite horizon goal-conditioned Markov Decision Process (MDP), denoted by the tuple $<\mathcal{G},\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},H>$ where $\mathcal{G}$ is the goal set, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the dynamic function, $\mathcal{R}$ is the reward function and $H$ is the horizon. At each timestep $t$, the mobile agent takes action $a_{t}\in\mathcal{A}$ (e.g., typing text, press button or slide the screen) based on its policy $\pi:\mathcal{S}\times\mathcal{G}\times\mathcal{A}\rightarrow[0,1]$, the current screenshot $s_{t}\in\mathcal{S}$, and a specific goal $g\in\mathcal{G}$ (e.g., search a new TV at Best Buy) selected in the beginning of each episode. The agent only receives the reward $r_{t}=1$ if the goal $g$ is accomplished, otherwise the reward $r_{t}=0$. The objective of the mobile agent is to find the policy $\pi$ which can accomplish all goals from the goal set $\mathcal{G}$ within the finite horizon $H$.
@@
For challenging control tasks with sparse and long-term reward signals, it is difficult to learn the useful policy arriving at the final goal within the finite horizon. Therefore, subgoal generation is particularly useful in providing the intermediate signals to facilitate learning. Then, we introduce the assumption of the existence of subgoals for the given goal, aiming to bring the goal-conditioned RL problem to the subgoal-conditioned RL problem as follows.
@@
Commonly adopted in literature (Sutton et al., 1999), the above assumption is mild and usually holds. For instance, when $N=1$, the subgoals and sub-trajectories are the original goal and trajectory, respectively. When $N=H$, each sub-trajectory is composed of one single transition-tuple $(s_{t},a_{t},r_{t},s_{t+1})$ with its corresponding subgoal.

### v2

# Advancing Autonomous VLM Agents via Variational Subgoal-Conditioned Reinforcement Learning

@@ Abstract
State-of-the-art (SOTA) reinforcement learning (RL) methods have enabled vision-language model (VLM) agents to learn from interaction with online environments without human supervision. However, these methods often struggle with learning inefficiencies when applied to complex, real-world decision-making tasks with sparse rewards and long-horizon dependencies. We propose a novel framework, Variational Subgoal-Conditioned Reinforcement Learning (VSC-RL), advancing the VLM agents in resolving challenging decision-making tasks. Fundamentally distinct from existing methods, VSC-RL reformulates the decision-making problem as a variational subgoal-conditioned RL problem with the newly derived optimization objective, Subgoal Evidence Lower BOund (SGC-ELBO), which comprises two key components: (a) maximizing the subgoal-conditioned return, and (b) minimizing the divergence from a reference goal-conditioned policy. We theoretically and empirically demonstrate that the VSC-RL can efficiently improve the learning efficiency without compromising performance guarantees. Across a diverse set of challenging benchmarks, including mobile device and web control tasks, VSC-RL consistently outperforms existing SOTA methods, achieving superior learning efficiency and performance.
@@ 1 Introduction
Recently, the large language models (LLMs) and vision-language models (VLMs) have demonstrated remarkable capabilities in content understanding and commonsense reasoning [37, 30, 11], achieving notable success in various real-world applications, such as visual question answering and visual captioning [5, 8]. These advancements highlight the strong potential of VLMs to tackle complex real-world decision-making problems (e.g., mobile device [34] and web control [45] tasks) via building intelligent VLM agents through the advanced VLMs [40, 44]. Meanwhile, after achieving impressive results in board games [31] and video games [6], reinforcement learning (RL) methods have been applied in online training VLM agents for tackling sequential decision-making tasks [4, 26].
@@
Overall, based on the specific training paradigm, VLM agents can be categorized into three main types: prompting-based, imitation-based, and RL-based agents. Directly leveraging VLMs (e.g., Gemini-1.5-Pro [33] and GPT-4V [21]) to capture the critical information from the multimodal content, prompting-based agents aim to generate action via prompting engineering and retrieving techniques [40, 38]. The performance of the prompting-based agents is usually limited, as the weights of these VLMs can not be updated. To address this limitation, some studies [42, 12] employ imitation learning techniques to fine-tune the open-source VLMs using human demonstrations. However, the performance of imitation-based agents is highly dependent on the quality and diversity of the demonstrations. Consequently, imitation-based agents may struggle with generalization and often underperform on out-of-distribution and unseen tasks. Recently, RL-based agents have emerged as a promising solution. By incorporating RL techniques, these agents enable VLMs to tackle complex sequential decision-making problems [4, 26]. Nevertheless, existing RL-based agents often suffer from the learning efficiency issue in addressing challenging control tasks with sparse reward signals and complicated goals. In many real-world scenarios, tasks require executing long sequences of actions, with rewards only provided upon successful completion. This delayed feedback poses a significant challenge for learning, fundamentally impacting the efficiency of RL-based agents. Some existing works attempt to address this issue by introducing implicit curriculum [3] or hand-crafted subgoals [10, 7]. However, these existing approaches often fail to learn a proper policy in the real-world complex sequential decision-making task due to the complicated subgoal generation and curriculum design methodologies.
@@
To address the fundamental limitations of RL-based agents mentioned above, we introduce Variational Subgoal-Conditioned RL (VSC-RL), a novel RL-based VLM agent method for enhancing learning efficiency in real-world complex sequential decision-making tasks. Based on the perspective of variational inference, VSC-RL reformulates the decision-making task as the variational subgoal-conditioned RL problem, which is later efficiently solved by utilising extensive optimization techniques. Additionally, VSC-RL utilises the significant reasoning and planning capabilities of VLM to autonomously decompose the complex goal into feasible subgoals. Given the generated subgoals, VSC-RL optimizes the objective of SubGoal-Conditioned Evidence Lower BOund (SGC-ELBO), thus effectively improving learning efficiency, consisting of (a) maximizing the subgoal-conditioned return of the target agent and (b) minimizing the subgoal-conditioned difference with the reference agent. We theoretically derive the new objective of SGC-ELBO from the original optimization objective, ensuring both improved efficiency and performance guarantees. Empirical results on various benchmarks validate our statement that VSC-RL significantly outperforms SOTA VLM agents in both sample efficiency and final performance.
@@
In this paper, literature related to VLM agents and RL methods is discussed in Section 2. We introduce notations related to goal-conditioned RL, variational RL and subgoal generator in Section 3. In Section 4, we illustrate how to formulate the sequential decision-making problem as a variational subgoal-conditioned RL problem and derive the new optimization objective: SGC-ELBO, followed by the practical implementation of VSC-RL. In Section 5, the experimental results over various benchmarks exhibit that our VSC-RL agent can achieve superior performance compared to existing SOTAs. Overall, the main contributions of this paper are summarised as follows:
@@
  • We propose VSC-RL, a novel variational subgoal-conditioned RL method for enhancing VLM agents in resolving real-world sequential decision-making problems.
@@
  • We theoretically show that SGC-ELBO, the optimization objective of the VSC-RL, can effectively improve learning efficiency while maintaining the performance guarantee.
@@
  • We experimentally show that VSC-RL significantly outperforms various SOTAs in both learning efficiency and final performance on various challenging benchmarks.
@@ 2 Related Works
2.1 VLM Agents for Decision-making
In real-world complex control tasks requiring capacities in reasoning, planning and content understanding, it is necessary to enable agents with the vision-language models (VLMs). In particular, the VLMs can process and abstract the image and language content for challenging decision-making tasks, especially in mobile device and web control tasks [34, 17]. Existing VLM agents can be categorized as prompting-based, imitation-based, and RL-based agents based on the corresponding learning paradigms. Additionally, some recent works explore using VLM to enhance agents’ abilities.
@@ 2.1 VLM Agents for Decision-making
Leveraging the inherent reasoning and planning abilities of prosperity VLMs (e.g., Gemini-1.5-Pro [33] and GPT-4V [21]), the prompting-based agent makes decision via prompting engineering and retrieving techniques. For instance, AppAgent [40] first introduces a unified prompting-based agent method to enable the vision-language model to directly interact with mobile applications by providing the prompts with details of actions. Set-of-Marks [38] proposes a new prompting method to enhance the visual grounding ability of VLM. However, the performance of these prompting-based agents is always sensitive to the prompts required to be manually and carefully designed. Therefore, it is challenging for the prompting-based agent to directly output the correct and desired actions to address real-world complex control problems.
@@
The imitation-based agent learns to mimic the expert behaviours by fine-tuning the policy on human demonstration. Recently, Android in the Wild (AitW) [27] collected large-scale datasets of mobile device control tasks, enabling agents to directly learn from human experience. AutoUI [42] and CogAgent [12] fine-tune the VLM-based policies with the AitW dataset, remarkably outperforming the prompting-based agent. In order to adapt the fine-tuned agent to the online environment, Filtered BC [23] introduces online imitation mechanisms to learn from successful online experiences. Unfortunately, these methods rely heavily on high-quality human demonstrations and often struggle to generalize to unseen tasks, limiting their application in diverse real-world scenarios.
@@
Different to prompting-based and imitation-based agents, the RL-based agent can autonomously optimize the policy through trial-and-error interactions with environments, without human supervision. DigiRL [4] introduces a unified offline-to-online RL framework that enables agents to learn directly from real-time interactions in dynamic environments, improving performance without the need for curated datasets. DistRL [35] builds an asynchronous distributed RL system, allowing training multiple agents in parallel across different environments, thus significantly enhancing scalability and convergence speed. WebRL [26] introduces a self-evolving online curriculum RL framework, enabling effective training of web agents through adaptive task generation in web control tasks. However, these RL-based agents still fundamentally suffer from the learning efficiency issue in challenging sequential decision-making tasks with sparse rewards and long horizons.
@@
Recent works have shown that VLM can enhance the RL method via its remarkable capacities of reasoning, planning, and content understanding. Recent works suggest adopting VLM in reward-shaping for RL. For instance, VLM-RMs [28] demonstrate that VLMs can serve as effective reward models for learning complex skills. VLM can also generate the subgoals to guide the learning process for autonomous driving [22] and robot [39] tasks. Nonetheless, it is still an open problem how to effectively integrate the VLM-generated subgoals into RL.
To mitigate the above issues, we present VSC-RL, which can autonomously decompose the goal into feasible subgoals by advanced VLM, and then efficiently resolve each subgoal from the principle of variational inference.
@@ 2.2 Goal-conditioned and Variational RL
Sequential decision-making tasks can be viewed as the goal-conditioned RL problem [16]. Based on the current state, the agent aims to find the optimal policy that guides progress toward the given goal for maximizing the return. Hindsight experience replay [3] introduces an implicit curriculum learning method to enhance learning efficiency and robustness. With the perspective of divide-and-conquer, some approaches suggest guiding the agent with subgoals as intermediate reward signals via imagination [7, 19] and tree-search [14, 24].
@@
The RL problem can be viewed as the variational inference problem [15], which can be resolved by utilising extensive optimization tools, thus effectively improving the learning efficiency. Applying the expectation-maximization algorithm in the actor-critic method in RL, VIP [20] presents a unified variational inference framework. MPO [1, 2] proposes a series of off-policy RL with entropy regulation in the manner of expectation-maximization. VDPO [36] and CVPO [18] apply the variational inference techniques in addressing the RL problem with delayed signals and safety constraints, respectively.
This paper aims to show how to formulate the control problem as a variational subgoal-conditioned RL problem from the perspective of variational inference, which allows us to resolve the complicated control task by utilising extensive optimization tools.
@@ 3 Preliminaries
We formulate the RL problem as the finite horizon goal-conditioned Markov Decision Process (MDP), denoted by the tuple $<\mathcal{G},\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},H>$ where $\mathcal{G}$ is the goal set, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the dynamic function, $\mathcal{R}$ is the reward function and $H$ is the horizon. At each timestep $t$, the agent takes action $a_{t}\in\mathcal{A}$ (e.g., typing text, press button or slide the screen) based on its policy $\pi:\mathcal{S}\times\mathcal{G}\times\mathcal{A}\rightarrow[0,1]$, the current screenshot $s_{t}\in\mathcal{S}$, and a specific goal $g\in\mathcal{G}$ (e.g., search a new TV at Best Buy) selected in the beginning of each episode. The agent only receives the reward $r_{t}=1$ if the goal $g$ is accomplished, otherwise the reward $r_{t}=0$. The objective of the agent is to find the policy $\pi$ which can accomplish all goals from the goal set $\mathcal{G}$ within the finite horizon $H$.
@@
For challenging control tasks with sparse and long-term reward signals, it is difficult to learn a useful policy that arrives at the final goal within a finite horizon. Therefore, subgoal generation is particularly useful in providing the intermediate signals to facilitate learning. Then, we introduce the assumption of the existence of subgoals for the given goal, aiming to bring the goal-conditioned RL problem to the subgoal-conditioned RL problem as follows.
@@
Commonly adopted in literature [32], the above assumption is mild and usually holds. For instance, when $N=1$, the subgoals and sub-trajectories are the original goal and trajectory, respectively. When $N=H$, each sub-trajectory is composed of one single transition-tuple $(s_{t},a_{t},r_{t},s_{t+1})$ with its corresponding subgoal.